The price elasticity of demand is –1.5, and the share of the tax borne by consumers is 0.60. What is the price elasticity of supply?

  • 5

  • 3.75

  • 2.25

  • 0.4

  • correct answer not provided
Answer #1

Option 3. 2.25

Explanation:

The tax burden on consumer = Elasticity of supply/ (Elasticity of demand + Elasticity of supply)

0.60 = Elasticity of supply/(1.5 + Elasticity of supply)

0.60 *(1.5) + 0.60*elasticity of supply = Elasticity of supply

or, 0.90 = Elasticity of supply - 0.60*elasticity of supply

or, 0.90 = 0.40 * elasticity of supply

or, Elasticity of supply = 0.90/0.40 = 2.25